### v1

# Electron-nucleus cross sections from transfer learning

Abstract
Transfer learning (TL) allows a deep neural network (DNN) trained on one type of data to be adapted for new problems with limited information. We propose to use the TL technique in physics. The DNN learns the physics of one process, and after fine-tuning, it makes predictions for related processes. We consider the DNNs, trained on inclusive electron-carbon scattering data, and show that after fine-tuning, they accurately predict cross sections for electron interactions with nuclear targets ranging from lithium to iron. The method works even when the DNN is fine-tuned on a small dataset.
Transfer learning (TL) is a widely studied phenomenon in psychology and education [1]. It refers to the ability of a person who has learned skills in one specific field to easily acquire skills needed in related areas of life. The concept of TL has been recognized in the artificial neural network community since 1976 [2]. Recently, Andrew Ng highlighted its significance for developing modern deep neural network (DNN) systems [3]. There are numerous successful applications of TL in deep learning [4, 5]. One example could be DNN, trained on non-medical data, which, after fine-tuning, is used to analyze medical photos and diagnose cancer.
This letter proves that the TL technique can be efficient in modeling physical systems. The information acquired by a network that has learned the physics of a specific process is transferred, using deep learning methods, to describe related processes. The only requirement in this framework is that the network must be capable of learning the basic universal features that describe the process. Only minimal information is needed to obtain a model for a related process.
We focus on the problem of modeling electron-nucleus scattering. Our motivation comes from studies of neutrinos’ fundamental properties. The next generation of neutrino-oscillation experiments DUNE [6] and Hyper-Kamiokande [7], aim to determine the value of the Dirac phase and shed light on the question of the dominance of matter in the Universe. Their success is contingent upon the accuracy of our estimates of neutrino cross sections for interactions with nuclear targets, which is required to translate the observed neutrino-antineutrino differences between the appearance-event rates to the oscillation probabilities. In particular, interactions of (anti)neutrinos with argon will be detected in DUNE, and neutrino-oxygen scattering (in water) is relevant for Hyper-Kamiokande.
A number of factors contribute to the complexity of the problem of neutrino interactions with nuclei in the few-GeV energy region, relevant for the oscillation studies [8]. The measured cross sections are averaged over broad fluxes, which greatly diminishes differences between different interaction channels at the inclusive level. Different interaction mechanisms yield the same final states, making a piecewise model development approach unsuitable. Only nuclear cross-sections are available with the desired precision, which makes a clear-cut identification of nuclear effects challenging. Moreover, the process of scattering involves multicomponent vector and axial currents, which are impossible to disentangle and determine based on neutrino data alone.
Recently, a strong consensus has emerged to make full use of the similarities of electron and neutrino interactions at the nuclear level [9], to greatly reduce the uncertainties of neutrino cross sections originating from nuclear effects, such as the description of the ground state properties and final-state interactions.
The scarcity of the electron-scattering measurements for argon and oxygen is currently a major obstacle [10, 11, 12]. On the other hand, a large body of data is already available for carbon [13]. In this paper, we employ the concept of the TL to demonstrate that DNNs trained on carbon data can accurately predict the cross sections for nuclear targets ranging from lithium to iron, provided that even scarcely distributed measurements span the relevant kinematic region. Although the problem we consider is specific, the proposed method is general and can be applied to other particle and nuclear physics studies.
Before we delve into the electron-nucleus scattering, let us first clarify the concept of the TL. A deep neural network is an algorithm that transforms an input vector (or tensor) into an output (scalar, vector, or tensor). It can be represented by a graph of connected neurons grouped in multiple layers. A single neuron transforms the input into a scalar. In our analysis, we consider the DNNs containing ten blocks [14], each containing a fully connected layer of units and batch normalization layer.
Figure 1: Double-differential cross section ${d^{2}\sigma}/{d\omega d\Omega}$ for inclusive electron scattering on lithium [15], oxygen [11], aluminum [16], calcium [17], and iron [18] for selected kinematics. The fits obtained for the proportion of training to test datasets 7:3. The red line denotes the DNNs predictions, and the green area denotes $1\sigma$ uncertainty. The results for DNN with all/two last layers fine-tuned are shown in the top/bottom row. The electron-carbon DNN predictions multiplied by factor A/12 are shown by blue-dashed lines corresponding to $1\sigma$ uncertainty denoted by a light blue area. The red (green) points represent the training (test) dataset.
One of the most intriguing general properties of deep learning is that neurons in the first layers of DNN extract basic features from the input, while the units in the last layers combine the simple features into more complex ones. This mechanism is known as representation learning [19]. For instance, the initial layers extract basic features such as edges and corners in the visual analysis of picture. The later layers combine these features to make decisions for abstract problems like content recognition in the picture of animal species.
Assuming that the ability to extract basic features is universal for a broad range of problems, a network trained on one data type can be fine-tuned to analyze other loosely related data. It can be understood as transferring knowledge gained from studying one problem to a different problem. One common approach involves replacing the last layers of a pre-trained network with a new one and training them with data from the new problem [20]. Another possibility is to fine-tune all parameters of the pre-trained network slightly.
These approaches offer several benefits, the most significant being the ability to train a new model using a small dataset, even if it only partially covers the domain of interest. TL’s potential is vast, opening up new application possibilities in various domains.
To prove that TL works effectively and can be utilized to model lepton-nucleus scattering, we consider the DNNs trained on the inclusive electron-carbon cross section data [14]. The models obtained there predict cross sections in the kinematic range covering quasielastic peak, resonance region, and deep inelastic scattering region. This analysis considers the bootstrap model [14], which showed very good extrapolation abilities. In this approach, we collected an ensemble of $50$ networks. The mean and variance over all network predictions in the ensemble give the model’s predictions and predictive uncertainty, respectively. Each network in the ensemble has the same structure.
Figure 2: Caption same as in Fig. 1 but results obtained for the proportion of training to test datasets 1:9.
We presume the network, which was trained using electron-carbon data, has assimilated the fundamental nuclear structure effects that apply to all nuclei. While the physics governing the nucleus remains similar across all targets, there are variations among different types of targets. These differences are likely to be accumulated and captured in the final layers of the network rather than in the initial layers. Furthermore, if the DNN contains substantial information about nuclear physics, there may be no need for extensive new measurements to fine-tune it.
The following framework is applied to investigate how TL works for electron-nucleus scattering: The pre-trained DNNs are retrained using new measurements in two distinct scenarios. Firstly, the parameters in all layers of the DNNs are optimized. In the second scenario, we only optimize the parameters in the final two layers (fine-tuning only the last layer was ineffective). Additionally, to train the DNN, we divide the data into training and test sets. Initially, the split is 7:3, but to gauge the effectiveness of a limited training dataset, we also consider a split of 1:9.
We consider measurements of the inclusive cross section for electron scattering off lithium (6Li), oxygen (16O), calcium (40Ca), aluminum (27Al), and iron (56Fe) for energies up to $8$ GeV. The data are taken from the database [21] and they are in the form of double differential cross sections $d^{2}\sigma/d\omega d\Omega$ depending on incoming electron energy (E), energy transfer ($\omega$), and scattering angle ($\theta$). We excluded low-energy data points from each target dataset as in [14]. We consider a separate ensemble of 50 DNNs that are fine-tuned for each process to describe new measurements. To adapt the pre-trained (on electron-carbon scattering data) DNN models to analyze the other target data, we multiply the pre-trained DNN output by the scale factor of A/12, where A corresponds to an atomic number of a given target, 12 is the atomic number of carbon. The DNN takes for the input the electron energy, transfer of energy, scattering angle, and two related variables $\cos\theta$ and four-momentum transfer $(Q^{2})$.
To fine-tune the model, we only needed to run the optimizer for several hundred epochs, ranging from 450 to 1,200, compared to over 40,000 epochs required for optimizing the pre-trained model. We consider the same statistical framework as in Ref. [14]. The loss is given by the $\chi^{2}$, including statistical and point-to-point systematic uncertainties. Additionally, we include a penalty term for each independent dataset that allows us to control the relative normalization parameters and account for systematic normalization uncertainties [22]. More details about the analysis can be found in the supplement.
The electron-carbon scattering data covers a considerable kinematic region that mostly overlaps with the kinematic range of data collected for targets other than carbon. However, the carbon scattering data’s kinematic domain does not cover some non-carbon measurements. One example is aluminum data at $\theta\approx 53^{\degree}$.
Our computational work began by randomly splitting new data into training and test datasets in a 7:3 proportion for each target. Note that the test data are not included in the model’s tuning process. We observed a perfect agreement between the predictions of fine-tuned DNNs and the data in this case, as shown in Fig. 1 and the supplementary material. The best agreement is achieved when parameters in all layers are fine-tuned, but the model with only the last two layers fine-tuned works well, too.
In all figures, we plot the pre-trained DNN model predictions multiplied by the factor of A/12 to illustrate how the fine-tuning procedure changes the pre-trained model predictions. There are well-known similarities between the nuclei of 16O and 12C. Therefore, in this case, simple re-scaling of the cross sections produces a reasonable agreement. For the other targets, it is necessary to fine-tune the model more significantly. Our DNN models predict the cross sections in quasielastic peak, dip, resonance, and the onset of deep inelastic regions. Different nuclear mechanisms contribute to each kinematic domain; hence, a simple scaling cannot work.
The power of TL is revealed once the information in the training dataset is limited. To illustrate this, we consider proportions of 1:9 between the training and test datasets. In Fig. 2, we show some remarkable examples, and for other comparisons, refer to the supplementary material. The fully fine-tuned model (the first scenario) can reproduce the cross-section’s shape and strength with noteworthy quality for almost all targets, even when there are few or no training points in a particular set of measurements. In the second scenario, where the two last layers are only trained, the DNN predictions also agree with the data with reasonable accuracy. However, the fact that the model with all layers optimized works better shows that the distribution of the nuclear structure information in the network is nontrivial. Both fine-tuning scenario models, all and the last two layers optimized, work well for oxygen, calcium, and iron. For lithium and aluminum, fine-tuning all layers leads to the model, which is significantly more accurate in the predictions than the scenario in which the last two layers are optimized. Interestingly, the second scenario model has a problem with correctly predicting the strengths of quasielastic, $\Delta$-resonance peaks, and dip in between – if the peaks are properly matched, the dip is not.
The success of the TL-based approach can be attributed to two primary sources. Firstly, DNN apparently can identify relationships in electron scattering data [13], for instance, those described by the superscaling hypothesis [23]. The DNN can recognize these relationships by analyzing electron-carbon data and using minimal input information from measurements of other targets. The direct comparison of both approaches, DNN and superscaling, needs a dedicated study and goes beyond the scope of this paper. Secondly, DNNs are known for their strong generalization abilities [24], although the exact mechanism behind this still needs to be well understood. This property is related to the fact that DNNs have a much larger number of parameters than the data constraints. Therefore, when the model is trained, its parameters are only partially constrained by the data, leading each parameter to contain a fraction of random noise, preventing overfitting and producing good model generalization ability.
Our study demonstrates that deep neural networks can effectively learn the physics of electron-carbon scattering. The network encodes fundamental properties of the nuclear medium, which can be leveraged through the TL techniques to model related processes, even with limited information. This method is rooted in the fundamental property of deep learning – DNNs can extract basic features from data and use them to build more complex abstract features.
Our paper focuses on a specific aspect of lepton-nucleus scattering physics, but the potential of the TL extends far beyond this application. It can be effectively utilized in various other domains within nuclear and particle physics and other scientific fields. In particular, our proposed technique holds great promise in accurately modeling neutrino-nucleus scattering. There are significant parallels between the interactions of electrons and neutrinos with nuclei, suggesting that a model trained on electron-nucleus scattering data can be leveraged, through the TL, to model neutrino-nucleus cross sections. Although beyond the scope of this paper, this approach will be a focal point of our future research. Such a novel methodology has the potential to significantly enhance our understanding of neutrino interactions with nuclei, which is crucial for next-generation neutrino long-baseline oscillation experiments.
Acknowledgments—This research is partly (K.M.G., A.M.A., J.T.S.) or fully (B.E.K., R.D.B., J.L.B, H.P.) supported by the National Science Centre under grant UMO-2021/41/B/ST2/ 02778.